# Theory of enhanced-by-coincidence neural information transmission

### v1

@@ III.1 The diffusion approximation for the encoding signal
To build an interpretable theory of the NERM, we will first adopt the diffusion approximation for the readout neuron afferent current $I(t)$ in Eq. (3) [42, 45, 17]. This approximation is based on the fact that for high enough input spike count $x_{i}(t)\delta$ in a time window $\delta$ the Poisson statistics of the count $n_{i}(t+\delta)-n_{i}(t)$ approaches a normal distribution with mean and variance equal to $x_{i}(t)\delta$. In the diffusion approximation, the input current $I(t)$ and the readout neuron membrane potential $V(t)$ in Eq. (2) satisfy (see Appendix B and [17]):
@@
where $\overline{\mu}^{(s)}$ and ${\sigma_{\rm V}}^{2}$ are the average of current means and variances $\mu^{(s)}_{i}$ and ${\sigma_{\rm V}}_{i}^{2}$ over encoding neurons $i$, respectively. For simplicity, in Eqs. (6) we consider a homogeneous gain vector $w_{i}=w$ $\forall i$ (see Appendix B for a generalization to arbitrary correlation matrices $\bar{C}$ and gain vectors $\bf w$). The time-shifted covariance of the current $I(t)$ takes the form (from Eq. (5)):
@@
where ${v_{\rm c}}$ is the stationary variance of the colored noise.
@@
In Appendix D we assess the validity limits of the diffusion approximation and its analogy with the non-Markovian random walk problem.
@@ III.2 The quenched-noise approximation for the LIF neuron rate mean
In the following, we derive expressions for the mean and variance (and SNR, see Eq. 4) of the readout neuron defined by Eq. 5 as a function of the parameters ${\mu_{\rm I}}^{(\pm)},{v_{\rm c}},{\tau_{\rm c}},w,{\tau_{\rm m}}$.
@@
Here we present an approximated expression for ${r_{\rm R}}$ in the ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ limit, in the quenched-noise approximation. The quenched-noise approximation generalizes the Moreno solution [17] to larger values of $\alpha_{\rm M}$. The approximation stands on a simple argument: whenever ${\tau_{\rm c}}\gg{\tau_{\rm m}}$, in time windows of length $\ell$, ${\tau_{\rm m}}\ll\ell\ll{\tau_{\rm c}}$, the coloured noise $\xi(t)$ is approximately constant and, consequently, the rate of the LIF neuron behaves as that of a white-noise LIF neuron with input current $I(t)=\tilde{\mu}+{\sigma_{\rm I}}\,\eta(t)$, where $\tilde{\mu}={\mu_{\rm I}}+{\sigma_{\rm c}}\,\xi$ is an effectively constant drift. Under this assumption, the firing rate ${r_{\rm R}}$ in time intervals of size ${T_{\rm c}}\gg{\tau_{\rm c}}$ can be approximated as the average of ${r_{\rm w}}(\tilde{\mu})$ over the stationary distribution of values of $\tilde{\mu}\sim{\cal N}(\tilde{\mu};{\mu_{\rm I}},{v_{\rm c}})$ with ${v_{\rm c}}={\sigma_{\rm c}}^{2}/(2{\tau_{\rm c}})$ (see (Eq. 5)), where ${\cal N}(\cdot;m,v)$ denotes the normal distribution with mean $m$ and variance $v$. Thus, the readout firing rate in the quenched-noise approximation ${r_{\rm R}}({\mu_{\rm I}})\simeq{r_{\rm q}}({\mu_{\rm I}})$ is given by:
[…]
where ${r_{\rm w}}(\mu)$ is the Siegert solution for the mean rate of a white noise LIF neuron (see Eq. (37) in Appendix E) [46, 42]. In the saddle-point approximation for large ${\tau_{\rm c}}/{\tau_{\rm m}}$ (see Appendix E) we have:
@@
Interestingly, if we now expand $[1-{v_{\rm c}}g^{\prime}_{1}({\mu_{\rm I}})]^{-1/2}$ in (Eq. 10) to the first order in $\nu_{\rm W}=({\sigma_{\rm c}}^{2}{\tau_{\rm m}})/(\sigma_{I}^{2}{\tau_{\rm c}})$ (W stands for Walter et al, who use the same perturbative parameter in reference [49]), and the exponential in (Eq. 10) to the first order in ${\tau_{\rm m}}/{\tau_{\rm c}}$, we recover the Moreno et al expression for ${r_{\rm R}}$ at first order in ${\tau_{\rm m}}/{\tau_{\rm c}}$ in Eq. (36). This is therefore a generalization of the Moreno solution for ${r_{\rm R}}$ in the ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ limit [17, 19], expected to hold for larger values of $\alpha_{\rm M}$ (Figs. 4,8-A and Appendix F).
@@
Comparing the Moreno [17] and quenched-noise approximations with the numerical data in the limit ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ reveals that: (i) for low ${\sigma_{\rm c}}$ both approximations $r_{\rm M}$ and ${r_{\rm q}}$ converge to the white-noise LIF activation function ${r_{\rm w}}$ (Eq. (10)) and for low enough ${\sigma_{\rm c}}$ and hence $\alpha_{\rm M}$, both solutions $r_{\rm M}$ and ${r_{\rm q}}$ are consistent with each other (Fig. 4-A red lines, Fig. 8-A and Appendix F); (ii) for larger values of ${\sigma_{\rm c}}$ the solutions differ, and the numerical estimation of ${r_{\rm R}}$ (i.e. $\tilde{r_{\rm R}}$; see Appendix C), is consistent only with the quenched-noise approximation (Fig. 4-B blue lines); (iii) to find full agreement between numerical simulations and ${r_{\rm q}}$, we need to apply the finite-Number of Integration Steps (NIS) correction (Fig. 4-B, see Appendix C).
@@
Figure 4: Readout firing rate ${r_{\rm R}}$ vs ${\mu_{\rm I}}$: numerical estimations and analytical approximations. (A) ${r_{\rm R}}$ vs ${\mu_{\rm I}}$ for two values of the noise correlation alignment: ${\alpha_{\rm V}}=0$ (no noise correlations, red curve), and ${\alpha_{\rm V}}=0.9$ (strong information-limiting noise correlations, blue curve). The numerical estimation (points) correspond to a simulation with the highest NIS, ${n_{\rm is}}=2^{17}$ (see Appendix C). The error-bars are the rate standard deviation in the unit time window, $(\tilde{v_{\rm R}}\times{T_{\rm c}})^{1/2}$, estimated as the standard deviation of the spike count ${n_{\rm R}}({T_{\rm c}})$ in $[0,{T_{\rm c}}]$ across ${n_{\rm tr}}=10^{3}$ simulation trials. The analytical activation functions corresponding to the white-noise LIF, ${r_{\rm w}}({\mu_{\rm I}})$ in (37), the Moreno solution $r_{\rm M}({\mu_{\rm I}})$ for ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ in Eq. (36) [17], and the quenched-noise approximation ${r_{\rm q}}$ in Eq. (9), are shown. The simulation parameters are: ${\tau_{\rm c}}=0.1$, ${\tau_{\rm m}}=0.005$, ${\sigma_{\rm V}}=1.778$, $w=0.450$, ${v_{\rm V}}=25.39$. (B) Numerical estimations of ${r_{\rm R}}$ vs ${\mu_{\rm I}}$ for ${\tau_{\rm c}}\simeq 0.720$, and all the values of the NIS, ${n_{\rm is}}=2^{m}$ for $m=12,\ldots,17$ (different green points at a common abscissa indicate different values of the NIS). The error-bars indicate the estimation $(\tilde{v_{\rm R}}{T_{\rm c}})^{1/2}$. We report as well the finite-NIS corrected estimation of $\tilde{r_{\rm R}}({\mu_{\rm I}})$ (blue diamonds) and their error in the regression (error-bars), as well as the analytical approximations $r_{\rm M}$ and ${r_{\rm q}}$. The finite-NIS corrected estimation is compatible with the quenched-noise approximation only. See Appendix C for details, and more on the comparison between the raw numerical, NIS-corrected numerical, Moreno, quenched-noise in Appendix F. The simulation parameters are: ${\tau_{\rm c}}=0.720$, ${\sigma_{\rm c}}=30.593$, ${\tau_{\rm m}}=0.005$, $w=0.45$, ${v_{\rm V}}=9.8$.
@@ III.3 Intuitive interpretation of the enhanced firing rate effect in the quenched-noise approximation
The quenched-noise approximation provides us with an intuitive interpretation of the enhancement of the firing rate in the presence of colored noise. Since ${r_{\rm q}}({\mu_{\rm I}})\simeq\left<{r_{\rm w}}({\tilde{\mu}})\right>_{\tilde{\mu}}$ (Eq. (9)), we expect the introduction of a colored noise to increase the LIF firing rate, ${r_{\rm R}}({\mu_{\rm I}})>{r_{\rm w}}({\mu_{\rm I}})$, for those values of ${\mu_{\rm I}}$ in which ${r_{\rm w}}$ is convex, and vice-versa – since $f(\mu)\right>_{\mu}>f(\left<\mu\right>_{\mu})$ is the definition of convex function. Precisely, since ${v_{\rm c}}^{1/2}$ is the standard deviation of ${\tilde{\mu}}$, we expect ${r_{\rm q}}({\mu_{\rm I}})>{r_{\rm w}}({\mu_{\rm I}})$ to hold for values of ${\mu_{\rm I}}$ and ${v_{\rm c}}$ such that ${\mu_{\rm I}}$ is lower, in units of ${v_{\rm c}}^{1/2}$, than the inflection point $\mu^{*}$ of the white-noise LIF neuron (satisfying ${\rm d}^{2}{r_{\rm w}}(\mu)/{\rm d}\mu^{2}|_{\mu^{*}}=0$). Fig. 5 shows that this is the case for both the quenched-noise and the Moreno approximations.
The enhancement in firing rate therefore has a simple explanation in the large-correlation timescale regime. The input current signal $I(t)$ fluctuates within a timescale ${\tau_{\rm c}}$. Although those fluctuations are symmetric around the mean ${\mu_{\rm I}}$, since the activation function of the white LIF model is non-linear, and convex for low enough current ${\mu_{\rm I}}$, the gain in firing rate for positive fluctuations, ${r_{\rm w}}({\mu_{\rm I}}+\delta)-{r_{\rm w}}({\mu_{\rm I}})$ is larger than the loss for negative fluctuations $-{r_{\rm w}}({\mu_{\rm I}}-\delta)+{r_{\rm w}}({\mu_{\rm I}})$. Thus, for the firing rate to be enhanced by colored noise, the LIF readout neuron must operate in a sufficiently low input current regime (so that ${r_{\rm w}}({\mu_{\rm I}})$ is convex, given the relevant parameters ${\sigma_{\rm I}},{\tau_{\rm m}}$, see Sec. V). This effect is known as Jensen’s force in the literature [50].

### v2

@@ III.1 The diffusion approximation for the encoding signal
To build an interpretable theory of the NERM, we will first adopt the diffusion approximation for the readout neuron afferent current $I(t)$ in Eq. (3) [42, 45]. This approximation is based on the fact that, for high enough input spike count $x_{i}(t)\delta$ in a time window $\delta$, the Poisson statistics of the count $n_{i}(t+\delta)-n_{i}(t)$ approaches a normal distribution with mean and variance equal to $x_{i}(t)\delta$. In the diffusion approximation, the input current $I(t)$ and the readout neuron membrane potential $V(t)$ in Eq. (2) satisfy (see Appendix B):
@@
where $\overline{\mu}^{(s)}$ and ${\sigma_{\rm V}}^{2}$ are the average of current means and variances $\mu^{(s)}_{i}$ and ${\sigma_{\rm V}}_{i}^{2}$ over encoding neurons $i$, respectively. For simplicity, in Eqs. (6) we consider a homogeneous gain vector $w_{i}=w$ $\forall i$ (see Appendix B for a generalization to arbitrary correlation matrices $\bar{C}$ and gain vectors $\bf w$). The time-shifted covariance of the current $I(t)$ takes the form (from Eq. (5b)):
@@
where ${v_{\rm c}}$ is the stationary variance of the colored noise. These results are consistent with those in reference [17].
@@
The NERM in the diffusion approximation defined by Eqs. (5) will be, from now on, our working hypothesis model, to which our analytical results refer, and to which the numerical simulations refer, unless stated differently. For an assessment of the validity of the diffusion approximation, and a study of the behavior of the SNR in the NERM model with spiking encoding current in Eqs. (1,2,3), we refer to the Appendix D.
@@ III.2 The quenched-noise approximation for the LIF neuron rate mean
In the following, we derive expressions for the mean and variance (hence, for the SNR, see Eq. (4)) of the readout neuron defined by Eq. (5) as a function of the parameters ${\mu_{\rm I}}^{(\pm)},{v_{\rm c}},{\tau_{\rm c}},w,{\tau_{\rm m}}$.
@@
Here we present an approximated expression for ${r_{\rm R}}$ in the ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ limit, in the quenched-noise approximation. The quenched-noise approximation generalizes the Moreno solution [17] to larger values of $\alpha_{\rm M}$. The approximation stands on a simple argument: whenever ${\tau_{\rm c}}\gg{\tau_{\rm m}}$, in time windows of length $\ell$, ${\tau_{\rm m}}\ll\ell\ll{\tau_{\rm c}}$, the coloured noise $\xi(t)$ is approximately constant and, consequently, the rate of the LIF neuron behaves as that of a white-noise LIF neuron with input current $I(t)=\tilde{\mu}+{\sigma_{\rm I}}\,\eta(t)$, where $\tilde{\mu}={\mu_{\rm I}}+{\sigma_{\rm c}}\,\xi$ is an effectively constant drift. Under this assumption, the firing rate ${r_{\rm R}}$ in time intervals of size ${T_{\rm c}}\gg{\tau_{\rm c}}$ can be approximated as the average of ${r_{\rm w}}(\tilde{\mu})$ over the stationary distribution of values of $\tilde{\mu}\sim{\cal N}(\tilde{\mu};{\mu_{\rm I}},{v_{\rm c}})$ with ${v_{\rm c}}={\sigma_{\rm c}}^{2}/(2{\tau_{\rm c}})$ (see (Eq. 5b)), where ${\cal N}(\cdot;m,v)$ denotes the normal distribution with mean $m$ and variance $v$. Thus, the readout firing rate in the quenched-noise approximation ${r_{\rm R}}({\mu_{\rm I}})\simeq{r_{\rm q}}({\mu_{\rm I}})$ is given by:
${r_{\rm q}}({\mu_{\rm I}})$ :=${r_{\rm w}}(\tilde{\mu})\right>_{{\cal N}(\tilde{\mu};\mu_{I},{v_{\rm c}})}\qquad\text{(quenched-noise approximation for ${\tau_{\rm c}}\gg{\tau_{\rm m}}$)}$ (9)
where ${r_{\rm w}}(\mu)$ is the Siegert solution for the mean rate of a white noise LIF neuron (see Eq. (42) in Appendix E) [46, 42]. In the saddle-point approximation for low enough ${v_{\rm c}}={\sigma_{\rm c}}^{2}/(2{\tau_{\rm c}})$, it can be proved (see Appendix E) that Eq. (9) takes the analytic form:
${r_{\rm q}}({\mu_{\rm I}})$ $={r_{\rm w}}({\mu_{\rm I}})\left[\frac{1}{1-{v_{\rm c}}g^{\prime}_{1}({\mu_{\rm I}})}\right]^{1/2}\exp\left(\frac{{v_{\rm c}}}{2}\frac{g_{1}^{2}({\mu_{\rm I}})}{1-{v_{\rm c}}g^{\prime}_{1}({\mu_{\rm I}})}\right),\qquad\text{${\tau_{\rm c}}\gg{\tau_{\rm m}}$, low ${v_{\rm c}}$}$ (10)
@@
Interestingly, if we now expand $[1-{v_{\rm c}}g^{\prime}_{1}({\mu_{\rm I}})]^{-1/2}$ in (Eq. 10) to the first order in $\nu_{\rm W}=({\sigma_{\rm c}}^{2}{\tau_{\rm m}})/(\sigma_{I}^{2}{\tau_{\rm c}})$ (W stands for Walter et al, who use the same perturbative parameter in reference [49]), and the exponential in (Eq. 10) to the first order in ${\tau_{\rm m}}/{\tau_{\rm c}}$, we recover the Moreno et al expression for ${r_{\rm R}}$ at first order in ${\tau_{\rm m}}/{\tau_{\rm c}}$ in Eq. (41). We conclude that Eq. (10) is therefore a generalization of the Moreno solution for ${r_{\rm R}}$ in the ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ limit [17, 19], expected to hold for larger values of $\alpha_{\rm M}$ (Figs. 5,9-A and Appendix G).
@@
Comparing the Moreno and quenched-noise approximations with the numerical data in the regime ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ reveals that: (i) for low ${\sigma_{\rm c}}$ both approximations $r_{\rm M}$ and ${r_{\rm q}}$ converge (see Eq. (10)) to the white-noise LIF activation function ${r_{\rm w}}$ and for low enough ${\sigma_{\rm c}}$ and hence $\alpha_{\rm M}$, both solutions $r_{\rm M}$ and ${r_{\rm q}}$ are consistent with each other (Fig. 5-A red lines, Fig. 9-A and Appendix G); (ii) for larger values of ${\sigma_{\rm c}}$ the solutions differ, and the numerical estimation of ${r_{\rm R}}$ (i.e. $\tilde{r_{\rm R}}$; see Appendix C), is consistent only with the quenched-noise approximation (Fig. 5-B blue lines); (iii) to find full agreement between numerical simulations and ${r_{\rm q}}$, we need to apply the finite-Number of Integration Steps (NIS) correction (Fig. 5-B, see Appendix C).
We stress that Eq. (10) is an analytical approximation of Eq. (9). As such, its validity limits are those of Eq. (9) (i.e., ${\tau_{\rm c}}\gg{\tau_{\rm m}}$), plus the further condition that ${v_{\rm c}}$ is low enough (see Appendix E). Throughout the article, with the exception of Appendix F, we always use our analytic expression for the rate mean, Eq. (10), and never the numerical evaluation of the integral in Eq. (9). As a matter of fact, our analytical expression for the rate mean is a good approximation, since the conditions for its derivation are satisfied in the range of parameters that we probe in our simulations. This is shown in Fig. 5-B and in Appendix E. The same holds for the expressions of the rate variance in the quenched-noise approximation, that we show below in Sec. III.4. Finally, we probe, in a designated simulation in Appendix F for constant ${\tau_{\rm c}}$ and increasing ${\sigma_{\rm c}}$ (hence, increasing ${v_{\rm c}}$), the quenched noise approximation beyond the validity limits of Eq. (10). Since, in this simulation, it is ${\tau_{\rm c}}\gg{\tau_{\rm m}}$, we are within the validity limits of the quenched-noise approximation and, indeed, the readout firing rate is still well explained, net of the finite NIS effect described below, by Eq. (9), although it is no longer well described by Eq. (10).
@@
Figure 5: Readout firing rate ${r_{\rm R}}$ vs input current ${\mu_{\rm I}}$: numerical estimations and analytical approximations. (A) ${r_{\rm R}}$ vs ${\mu_{\rm I}}$ for two values of the noise correlation alignment: ${\alpha_{\rm V}}=0$ (no noise correlations, red curve), and ${\alpha_{\rm V}}=0.9$ (strong information-limiting noise correlations, blue curve –see the value of ${\sigma_{\rm c}}$ in the legend). The numerical estimation (points) correspond to a simulation with the highest NIS, ${n_{\rm is}}=2^{17}$ (see Appendix C). The error-bars are the rate standard deviation in the unit time window, $(\tilde{v_{\rm R}}\times{T_{\rm c}})^{1/2}$, estimated as the standard deviation of the spike count ${n_{\rm R}}({T_{\rm c}})$ in $[0,{T_{\rm c}}]$ across the ${n_{\rm tr}}=10^{3}$ simulation trials. The analytical activation functions correspond to the white-noise LIF ${r_{\rm w}}({\mu_{\rm I}})$ in Eq. (42) (red line), the Moreno solution $r_{\rm M}({\mu_{\rm I}})$ for ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ in Eq. (41) [17] (dotted red and blue lines), and the quenched-noise approximation ${r_{\rm q}}$ in Eq. (9) (dashed red and blue lines). The simulation parameters are: ${\tau_{\rm c}}=0.1$, ${\tau_{\rm m}}=0.005$, ${\sigma_{\rm V}}=1.778$, ${N}=20$, $w=0.450$. (B) Numerical estimations of ${r_{\rm R}}$ vs ${\mu_{\rm I}}$ for ${\tau_{\rm c}}\simeq 0.720$, and all the values of the NIS, ${n_{\rm is}}=2^{m}$ for $m=12,\ldots,17$ (different green points at a common abscissa indicate different values of the NIS). We report as well the finite-NIS corrected estimation of $\tilde{r_{\rm R}}({\mu_{\rm I}})$ (blue diamonds) and their error in the regression (error-bars), as well as the analytical approximations $r_{\rm M}$ and ${r_{\rm q}}$. The finite-NIS corrected estimation is compatible with the quenched-noise approximation only. See Appendix C for details, and more on the comparison between the raw numerical, NIS-corrected numerical, Moreno, quenched-noise in Appendix G. Finally, we present a comparison of the analytical expression for the quenched-noise approximation for the mean rate, Eq. (10) with the general expression, Eq. (9). The simulation parameters are: ${\tau_{\rm c}}=0.720$, ${\sigma_{\rm c}}=30.593$, ${\tau_{\rm m}}=0.005$, $w=0.45$.
@@ III.3 Intuitive interpretation of the enhanced firing rate effect in the quenched-noise approximation
The quenched-noise approximation provides an intuitive interpretation of the enhancement of the firing rate in the presence of colored noise. Since ${r_{\rm q}}({\mu_{\rm I}})\simeq\left<{r_{\rm w}}({\tilde{\mu}})\right>_{\tilde{\mu}}$ (Eq. (9)), we expect the introduction of a colored noise to increase the LIF firing rate, ${r_{\rm R}}({\mu_{\rm I}})>{r_{\rm w}}({\mu_{\rm I}})$, for those values of ${\mu_{\rm I}}$ in which ${r_{\rm w}}$ is convex, and vice-versa – since $f(\mu)\right>_{\mu}>f(\left<\mu\right>_{\mu})$ is the definition of convex function. Precisely, since ${v_{\rm c}}^{1/2}$ is the standard deviation of ${\tilde{\mu}}$, we expect ${r_{\rm q}}({\mu_{\rm I}})>{r_{\rm w}}({\mu_{\rm I}})$ to hold for values of ${\mu_{\rm I}}$ and ${v_{\rm c}}$ such that ${\mu_{\rm I}}$ is lower, in units of ${v_{\rm c}}^{1/2}$, than the inflection point $\mu^{*}$ of the white-noise LIF neuron (satisfying ${\rm d}^{2}{r_{\rm w}}(\mu)/{\rm d}\mu^{2}|_{\mu^{*}}=0$). Fig. 6 shows that this is the case for both the quenched-noise and the Moreno approximations.
The enhancement in firing rate therefore has a simple explanation in the large-correlation timescale regime. The input current signal $I(t)$ fluctuates within a timescale ${\tau_{\rm c}}$. Although those fluctuations are symmetric around the mean ${\mu_{\rm I}}$, since the activation function of the white LIF model is non-linear, and convex for low enough current ${\mu_{\rm I}}$, the gain in firing rate for positive fluctuations, ${r_{\rm w}}({\mu_{\rm I}}+\delta)-{r_{\rm w}}({\mu_{\rm I}})$ is larger than the loss for negative fluctuations $-{r_{\rm w}}({\mu_{\rm I}}-\delta)+{r_{\rm w}}({\mu_{\rm I}})$. Thus, for the firing rate to be enhanced by colored noise, the LIF readout neuron must operate in a sufficiently low input current regime (so that ${r_{\rm w}}({\mu_{\rm I}})$ is convex, given the relevant parameters ${\sigma_{\rm I}},{\tau_{\rm m}}$, see Sec. V). This effect is known as Jensen’s force [50].